OpenStudy (anonymous):

f(x)=2+3*cos((π/10)(x−3)) Find the first positive value of x for which f(x) = 4 huh?

OpenStudy (jdoe0001):

$$ f(x)=2+3*cos\pmatrix{\cfrac{\pi}{10}(x−3)} \large \ ? $$

OpenStudy (anonymous):

i got 4.85

OpenStudy (jdoe0001):

hmm, well how did you get that?

OpenStudy (anonymous):

i put 4 in for x

OpenStudy (jdoe0001):

"for which f(x) = 4" ^ "y"

OpenStudy (jdoe0001):

$$ f(x)=2+3*cos\pmatrix{\cfrac{\pi}{10}(x−3)}\\ 4=2+3*cos\pmatrix{\cfrac{\pi}{10}(x−3)}\\ \cfrac{2}{3} = cos\pmatrix{\cfrac{\pi}{10}(x−3)}\\ cos^{-1}\pmatrix{\cfrac{2}{3}} = cos^{-1}\pmatrix{cos\pmatrix{\cfrac{\pi}{10}(x−3)}}\\ \cfrac{2}{3} = \cfrac{\pi}{10}(x−3) $$

OpenStudy (jdoe0001):

well, I got a bit off hehe

OpenStudy (jdoe0001):

$$ f(x)=2+3*cos\pmatrix{\cfrac{\pi}{10}(x−3)}\\ 4=2+3*cos\pmatrix{\cfrac{\pi}{10}(x−3)}\\ \cfrac{2}{3} = cos\pmatrix{\cfrac{\pi}{10}(x−3)}\\ cos^{-1}\pmatrix{\cfrac{2}{3}} = cos^{-1}\pmatrix{cos\pmatrix{\cfrac{\pi}{10}(x−3)}}\\ 48.19 = \cfrac{\pi}{10}(x−3) $$

OpenStudy (jdoe0001):

then you'd just need to solve for "x"

OpenStudy (whpalmer4):

you're trying to solve \[4 = 2+3\cos((\frac{\pi}{10})(x-3))\]Subtract 2 from both sides to give\[2=3\cos((\frac{\pi}{10})(x-3))\]Divide both sides by 3 to give\[\frac{2}{3}=\cos((\frac{\pi}{10})(x-3))\]Find the first value of \(\theta\) where \(\cos(\theta) = 2/3\) Then solve \(\theta = \pi/10(x-3)\) for \(x\)

OpenStudy (anonymous):

i got .3141?

OpenStudy (jdoe0001):

I got a much bigger degree number

OpenStudy (whpalmer4):

well, degrees = 57.3 * radians

OpenStudy (whpalmer4):

I get a number close to .3141, but I'm suspicious that you've just got pi/10 there :-)

OpenStudy (jdoe0001):

ahemm, 0.3 radians is barely 18 degrees, I got a much bigger value

OpenStudy (jdoe0001):

$$ \text{rounding off 48.19 to 48}\\ 48 = \cfrac{\pi}{10}(x−3)\\ \cfrac{48\times 10}{\pi} = x-3 $$

OpenStudy (whpalmer4):

you have to do the problem in radians...

OpenStudy (jdoe0001):

hmm, I see, "4" as in 4 radians, :), well, just the 48 will change

OpenStudy (anonymous):

so do youchange 48 to radians then solve?

OpenStudy (whpalmer4):

This approach will get you a solution but it won't be the first positive value of x that makes it true.

OpenStudy (whpalmer4):

Here's a plot of y=4 vs y=2+3cos[(pi/10)*(x-3)]